OpenStudy (anonymous):

~Medal+Fanning~ For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 9 3x + 4y = 26

OpenStudy (wolf1728):

You have a job weighing trucks. One day, 2 trucks (type x) and 1 truck (type y) are weighed together and their total weight is 9 tons. Another day you weigh 3 trucks (type x) with 4 (type y) trucks and find their total weight is 26 tons. What does each truck weigh? A) 2x + y = 9 B) 3x + 4y = 26 Multiply equation A by -4 A) -8x -4y = -36 B) 3x + 4y = 26 Adding the equations -5x = -10 x = 2 tons Putting the 2 into equation B 6 + 4y = 26 4y = 20 y = 5 tons

OpenStudy (anonymous):

This is all part 1?

OpenStudy (anonymous):

No wait this is part a and part b!! You're really awesome